A number is 12 less than its square. Find all such numbers.

Answer by richwmiller(9144)

x=x^2-12
0=x^2-x-12
find factors of -12 which add up to -1
-4 and 3
x^2-4x+3x-12=0
x*(x-4)+3*(x-4)=0
(x+3)*(x-4)=0
x=-3
x=4